OpenStudy (anonymous):

wrtie the standard form of the question of each line given the slope and y-intercept. slope=-3/5, y-intercept=5

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

y = (-3/5)x + 5

OpenStudy (anonymous):

what formular you use?

OpenStudy (anonymous):

this is the slope intercept form right? - i'm not from the us so i'm not sure of the terms and you require the standard form?

OpenStudy (anonymous):

the formula i used is y = mx + c

OpenStudy (anonymous):

3x+5y=25

OpenStudy (anonymous):

m =slope

OpenStudy (anonymous):

m=(y1-y2)/(x1-x2)

OpenStudy (anonymous):

opp

OpenStudy (anonymous):

right thats standard form you get this by multiplying by 5 so 5y = -3x + 25 3x + 5y = 25 - yes

OpenStudy (anonymous):

am wrong y2-y1 divide x2-x1

OpenStudy (anonymous):

yes m=(y1-y2)/(x1-x2) is correct formula for slope

OpenStudy (anonymous):

how you find y-intercept?

OpenStudy (anonymous):

well it depends on which points you call y1 , y2 etc

OpenStudy (anonymous):

to find y intercept put x = 0 in ter equation and work out value of y

OpenStudy (anonymous):

ok I understand

OpenStudy (anonymous):

you have pretty good knowledge

OpenStudy (anonymous):

you better than me

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw